OpenStudy (anonymous):

7a+2b+5+5a+9b+1 please help me. I am very new to this

OpenStudy (freckles):

What do you want to do?

OpenStudy (freckles):

Combine like terms?

OpenStudy (anonymous):

I'm new at this. I don't know how to do it. I need help understanding what to do

OpenStudy (freckles):

Can you tell me what the instructions are?

OpenStudy (anonymous):

The instructions says, simplify. That's all

OpenStudy (freckles):

Ok so we need to add to combine like terms

OpenStudy (freckles):

So like terms have the same variable part

OpenStudy (freckles):

like 7a and 5a are like terms

OpenStudy (freckles):

so they can be combined

OpenStudy (freckles):

the operation between 7a and 5a is addition what is 7+5?

OpenStudy (anonymous):

12

OpenStudy (freckles):

so 7a+5a=12a

OpenStudy (freckles):

do you notice any other like terms in your expression above?

OpenStudy (anonymous):

2b+9b

OpenStudy (freckles):

and what is 2b+9b?

OpenStudy (anonymous):

11 b

OpenStudy (freckles):

yes you also have another pair of like terms 5 and 1 the operation between them is also addition and 5+1=

OpenStudy (anonymous):

6

OpenStudy (freckles):

right we are done we just need to put everything together

OpenStudy (freckles):

we said 7a+5a=12a and 2b+9b=11b and 5+1=6 so we have 12a+11b+6

OpenStudy (anonymous):

ok

OpenStudy (freckles):

7a+2b+5+5a+9b+1 7a+5a+2b+9b+5+1 (7a+5a)+(2b+9b)+(5+1) (12a)+(11b)+(6) 12a+11b+6

OpenStudy (freckles):

any questions?

OpenStudy (anonymous):

Yes so is12a+11b+6 the answer?

OpenStudy (freckles):

Let me ask you if you think it is the answer. Do you see any other like terms to combine?

OpenStudy (anonymous):

Or do I continue to add ? No I dont

OpenStudy (anonymous):

I don't see any like terms

OpenStudy (anonymous):

Thank you very much!